OpenStudy (anonymous):

graph the inequality x+y<=4

OpenStudy (anonymous):

subtract x from both sides

OpenStudy (anonymous):

shade the part below the line

OpenStudy (anonymous):

what does "shade the part below the line " mean

OpenStudy (raden):

first, graph the line of x+y=4 by using intercepts if x = 0 then y = 4, so get y intercept at point (0,4) if y = 0 then x = 4, so get x intercept at point (4,0) now conecting both point to make a line : 

OpenStudy (raden):

now we have to shading the solution area of x+y <= 4 (it is underline of line x+y=4)